OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y(less than or = to) x - 5 y(less than or = to) -x - 4 1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (anonymous):

Help please :) lol

OpenStudy (anonymous):

let's take one by one ! -5-5=-10 > 2 false so it can't be the right answer 5-5=0>-2 true...-5-4=-9>-2 False it can't be the correct answer

OpenStudy (anonymous):

all the answer's are false !

OpenStudy (anonymous):

sorry look at it again...im sorry i did that wrong..! ;\

OpenStudy (anonymous):

okk...did you understund the way how we solve this questions ?!

OpenStudy (anonymous):

i think so the answer would be A

OpenStudy (anonymous):

(x,y)------>(1,10) ........ 10 is y and x is 1 is 10 less than(or equal) 1-5=-4 ?!